OpenStudy (frogbull2000):

please help!!! how do I convert y = -3x^2 + 3x +6 to the vertex form?

OpenStudy (frogbull2000):

standard form is y = ax^2 + bx + c and I need to get it to vertex form which is y - k = a(x - h)^2

OpenStudy (xapproachesinfinity):

complete the square! -3x^2+3x+6=-3(x^2-x)+6 now complete the procedure